translate each sentence into an equation. If four times the sum of a number and 5 is divided by 3, the result is 16.
Answer by jojo14344(1513)

Let's assigned "x" to the unknown number.
So,
4 (x+5)/3 = 16
x+5 = 48/4
x = 12-5
x = 7 = unknown number
To check,
4 (7+5)/3 = 16
4 (12/3) = 16
4 * 4 = 16
16=16